# ViEEG: Hierarchical Neural Coding with Cross-Modal Progressive Enhancement for EEG-Based Visual Decoding

### v2

@@ Abstract.
Understanding and decoding brain activity into visual representations is a fundamental challenge at the intersection of neuroscience and artificial intelligence. While EEG-based visual decoding has shown promise due to its non-invasive, low-cost nature and millisecond-level temporal resolution, existing methods are limited by their reliance on flat neural representations that overlook the brain’s inherent visual hierarchy. In this paper, we introduce ViEEG, a biologically inspired hierarchical EEG decoding framework that aligns with the Hubel-Wiesel theory of visual processing. ViEEG decomposes each visual stimulus into three biologically aligned components-contour, foreground object, and contextual scene-serving as anchors for a three-stream EEG encoder. These EEG features are progressively integrated via cross-attention routing, simulating cortical information flow from V1 to IT to the association cortex. We further adopt hierarchical contrastive learning to align EEG representations with CLIP embeddings, enabling zero-shot object recognition. Extensive experiments on the THINGS-EEG dataset demonstrate that ViEEG achieves state-of-the-art performance, with 40.9% Top-1 accuracy in subject-dependent and 22.9% Top-1 accuracy in cross-subject settings, surpassing existing methods by over 45%. Our framework not only advances the performance frontier but also sets a new paradigm for biologically grounded brain decoding in AI.
Neuroscience Applications, Cognitive Science, Neural Decoding, Multimodal Learning, Zero-shot Learning, Image Reconstruction
copyright: acmlicensedjournalyear: 2025doi: XXXXXXX.XXXXXXXconference: Make sure to enter the correct conference title from your rights confirmation email; Oct 27–31, 2025; Dublin, Irelandisbn: 978-1-4503-XXXX-X/2018/06ccs: Human-centered computing HCI theory, concepts and modelsccs: Computing methodologies Hierarchical representationsccs: Computing methodologies Object recognitionccs: Computing methodologies Reconstruction
Figure 1. EEG brain visual decoding of ViEEG.
[…]
1. Introduction
@@
Electroencephalogram (EEG)-based visual decoding aims to bridge the gap between cortical dynamics and machine perception by reconstructing visual experiences from brain activity (Du et al., 2023). As a pivotal modality in brain-computer interfaces, EEG (Craik et al., 2019) has gained prominence due to its cost-effectiveness, portability, and millisecond-level temporal resolution. Recent years have witnessed substantial progress in EEG-based visual decoding methodologies (Du et al., 2023; Song et al., 2024; Liu et al., 2024; Bai et al., 2023; Fu et al., 2025; Wei et al., 2024; Li et al., 2024; Zhang et al., 2024). However, unlike functional magnetic resonance imaging (fMRI) signals (Worsley et al., 2002) that have established sophisticated decoding frameworks (Zafar et al., 2015; Horikawa and Kamitani, 2017; Rakhimberdina et al., 2021; Allen et al., 2022; Scotti et al., 2023; Sun et al., 2023; Huo et al., 2024; Wang et al., 2024; Zhou et al., 2024), EEG-based visual decoding remains in its infancy, with significant challenges unaddressed, particularly in modeling the hierarchical visual processing mechanisms of the brain.
Figure 2. Tri-View Hierarchical Framework for EEG-Based Visual Decoding.
@@
Contemporary EEG decoding approaches predominantly establish direct visual-semantic mappings between EEG signals and stimulus images. For instance, Wei et al. (Wei et al., 2024) employ contrastive learning for direct EEG-image embedding alignment with contrastive language-image pretraining (CLIP) (Radford et al., 2021; Ramesh et al., 2022), while Song et al. (Song et al., 2024) develop spatiotemporal neural architectures tailored for EEG visual decoding. Recent innovations include incorporating brain topological connectivity (Li et al., 2024) and integrating multimodal semantic cues (Zhang et al., 2024). Despite these advancements, most existing methods overlook a fundamental property of the human visual system: its hierarchical nature. We term this issue as Hierarchical Neural Encoding Neglect (HNEN). Rooted in the Hubel-Wiesel model theory (Ben-Yishai et al., 1995), which established the canonical hierarchical cortical processing pathway, the human visual system processes stimuli through dissociable stages: from low-level edge detection to high-level semantic integration (Felleman and Van Essen, 1991; Riesenhuber and Poggio, 1999). In contrast, existing methods collapse this biological hierarchy into flat representation learning, resulting in incomplete visual information capturing.
@@
Substantial neuroscientific evidencee (Ben-Yishai et al., 1995; Felleman and Van Essen, 1991; Riesenhuber and Poggio, 1999; Li et al., 2022) supports a hierarchically organized cortical visual pathway for visual perception: The early visual cortex (V1111Primary visual cortex (V1): Edge processing/V2222Secondary visual cortex (V2): Edge processing (Bridge et al., 2005)) specializes in elementary low-level feature extraction (e.g., edge contours), the intermediate ventral stream (V4333Visual Area 4 (V4): Shape/color processing/IT [^4] Inferotemporal Cortex (IT): Object recognition (Hansen et al., 2007; Arcaro and Livingstone, 2021)) processes object-level semantics, while higher-order association cortex correlates555association cortex correlates (ACC): Scene context integration (Grill-Spector and Malach, 2004) integrate high-level contextual scene understanding. This hierarchical visual decoding enables humans to progressively distill visual information—from coarse contours to fine-grained semantics—through layered cortical interactions. However, existing EEG-based brain decoding methods adopt a monolithic processing paradigm that adopts a flat representation paradigm that collapses biological hierarchy. By attempting direct mappings between EEG signals and holistic image embeddings, these approaches suffer from HNEN, failing to distinguish:
@@
  • Contextual scene attributes (association cortex correlates)
The absence of cortically-aligned feature disentanglement fundamentally limits their capacity to localize visual saliency, particularly in preserving contour fidelity during EEG decoding.
@@
  • A biologically-inspired hierarchical framework for EEG-based visual decoding, pioneering tri-stream modeling (Contour $\to$ Object $\to$ Context) aligned with cortical processing stages. To our knowledge, this is the first work to explicitly enforce feature disentanglement along the cortical hierarchy.
@@
  • New state-of-the-art performance on THINGS-EEG: achieving 40.9% Top-1 (↑50%) in subject-dependent and 22.9% Top-1 (↑45%) in cross-subject settings, substantially outperforming previous benchmarks.
2. Related Works
2.1. Multimodal Learning
Human perception operates through intrinsically multimodal integration, where the brain dynamically synthesizes information across sensory and cognitive domains. This biological foundation has inspired computational frameworks that align neural activity with perceptual modalities such as vision and language. Recent advances attempt to decode brain signals into interpretable outputs. Recent advances have demonstrated the feasibility of translating brain activity into interpretable outputs, including neural speech synthesis (Chen et al., 2024), EEG-to-video generation (Liu et al., 2024), and EEG-based image reconstruction (Song et al., 2024).
A dominant line of research leverages CLIP’s unified embedding space for cross-modal alignment. Mapping EEG signals into the CLIP space using InfoNCE-based contrastive learning has emerged as a standard strategy (Song et al., 2024; Wei et al., 2024; Li et al., 2024). However, these approaches typically treat EEG as a holistic signal, overlooking the brain’s hierarchical structure in visual processing. In contrast, our method ViEEG incorporates biologically grounded visual decompositions—such as edge-based contours and background-free object entities—to preserve both low-level structure and high-level semantics. On the THINGS-EEG dataset, ViEEG achieves 40.9% Top-1 accuracy, surpassing CLIP-aligned flat decoders by nearly 50%, establishing a new biologically informed paradigm with improved interpretability and efficiency.
2.2. Brain Visual Decoding
@@
Decoding the brain’s response to visual stimuli has been a central pursuit in neuroscience (Ben-Yishai et al., 1995; Felleman and Van Essen, 1991; Riesenhuber and Poggio, 1999; Li et al., 2022) and brain-computer interface research (Lee et al., 2020; Shi et al., 2024). Early efforts predominantly focused on fMRI-based visual decoding (Allen et al., 2022; Scotti et al., 2023; Sun et al., 2023; Huo et al., 2024; Wang et al., 2024), where high spatial resolution inherently captures hierarchical visual processing. For instance, Scotti et al. proposed MindEye (Scotti et al., 2023), mapping fMRI signals to CLIP features via diffusion models, while Wang et al. developed MindBridge (Wang et al., 2024) to align fMRI with images and textual captions. These methods leverage fMRI’s spatial precision to distinguish low-level (e.g., edges) and high-level (e.g., semantics) visual features implicitly, even when relying on simplistic fMRI-CLIP alignment.
@@
In contrast, the spatial ambiguity of the EEG centimeter scale forces existing methods (Rakhimberdina et al., 2021; Du et al., 2023; Song et al., 2024; Wei et al., 2024; Li et al., 2024; Zhang et al., 2024) into flat representation paradigms that confuse hierarchical visual processing. As the dominant brain decoding approach of fMRI and EEG, direct EEG-image alignment fails to model the brain’s progressive refinement from edges to semantics, a limitation we term HNEN. While some attempts incorporate electrode topology (Li et al., 2024) or depth cues (Zhang et al., 2024), these merely address symptoms (e.g., spatial noise) rather than HNEN’s root cause: EEG’s inability to functionally isolate cortical processing stages.
This divergence highlights a critical gap: fMRI passively inherits hierarchical decoding through anatomy, whereas EEG demands active hierarchy simulation. Our proposed ViEEG bridges this gap via artificial cortical anchors—decomposing images into low-level edge masks (V1/V2 proxies), background-free entities (IT-level), and high-level contextual scenes (association cortex). By enforcing hierarchical EEG embedding disentanglement, we circumvent EEG’s spatial limitations while aligning with biological visual processing—a novel paradigm transcending both fMRI’s anatomy-dependence and EEG’s flat representations.
3. Preliminaries
In this section, we introduce the key notations used throughout the paper and formalize the task of EEG-based visual decoding.
3.1. Key Symbols
@@
$E$ $\mathbb{R}^{C\times T}$ Raw EEG signals
$I$ $\mathbb{R}^{H\times W}$ Raw image
$I_{b}$ $\mathbb{R}^{D\times D}$ Binary object mask image
$I_{f}$ $\mathbb{R}^{D\times D}$ Foreground object image
$I_{r}$ $\mathbb{R}^{D\times D}$ Raw scene image
$F_{b},F_{f},F_{r}$ $\mathbb{R}^{d}$ Contour/object/context EEG embedding
$C_{b},C_{f},C_{r}$ $\mathbb{R}^{d}$ Contour/object/context CLIP embedding
@@
Table 1. Key symbol definitions with biological correspondences. $C$: EEG channels, $T$: timepoints, $H\times W$: image resolution, $D\times D$: reshaped image resolution (512), $d$: embedding dimension (1024).
3.2. Task Formulation
@@
We formulate EEG-based visual decoding as a hierarchical zero-shot object recognition task under an open-world assumption, where test categories are disjoint from the training set. Given EEG recordings $E$ with $N$ samples elicited by visual stimuli $\{I_{b},I_{f},I_{r}\}$, our objective is to learn hierarchical EEG embeddings $\{F_{b},F_{f},F_{r}\}$ that align with the corresponding CLIP embeddings $\{C_{b},C_{f},C_{r}\}$ through hierarchical visual decoding. The learning objective for zero-shot object recognition is formulated as:
(1) $\min_{\theta}\frac{1}{N}\sum_{i=1}^{N}D(F^{i},C^{i}),$
@@
where $D(\cdot,\cdot)$ measures feature similarity. On the THINGS-EEG dataset, we train with class-disjoint stimuli and test on unseen categories. At inference time, the concatenated EEG embeddings ${F_{b},F_{f},F_{r}}$ are compared to CLIP embeddings using cosine similarity within an InfoNCE-style contrastive learning framework (Radford et al., 2021).
Figure 3. Overview of the ViEEG framework. The input image is decomposed into three biologically inspired views: binary mask, foreground object, and raw scene. Corresponding EEG responses are encoded by parallel spatiotemporal encoders, integrated via hierarchical cross-attention, and aligned with CLIP embeddings through hierarchical contrastive learning.
4. Method
4.1. Overall Architecture
@@
ViEEG introduces a biologically inspired hierarchical architecture for EEG-based visual decoding, simulating layered cortical processing. As illustrated in Figure 2, ViEEG comprises three synergistic components: (1) Hierarchical image decomposition, (2) Hierarchical EEG encoding, and (3) Hierarchical Contrastive Learning. During initialization, we extract CLIP embeddings from the three decomposed visual representations. In the training phase, Hierarchical EEG Encoding is employed to extract EEG features, which are then aligned with the image CLIP embeddings using Hierarchical Contrastive Learning. During testing, EEG embeddings extracted from the test set are matched to target samples based on similarity measurements for model evaluation.
4.2. Hierarchical Image Decomposition
@@
Inspired by the hierarchical structure of the human visual system, as described in the classical Hubel-Wiesel model, we decompose each stimulus image $I$ into three biologically aligned representations via semantic segmentation.
@@
Step 1: Image Processing. We employ the pre-trained BiRefNet [^6] https://huggingface.co/ZhengPeng7/BiRefNet (zheng2024bilateral), a state-of-the-art high-resolution segmentation model, to generate hierarchical visual representations. The raw image $I$ is first normalized to yield $I_{r}$, ensuring consistent dimensions and intensity scaling. Saliency detection is then applied, and a binary contour representation is obtained via thresholding:
(2) $I_{b}=\mathbb{I}\left(\mathcal{F}_{\text{BiRefNet}}(I_{r})>\tau\right)$
@@
where $\mathcal{F}_{\text{BiRefNet}}$ denotes the pre-trained BiRefNet model, $\mathbb{I}(\cdot)$ is the indicator function, and $\tau$ is threshold value. The resulting $I_{b}$ serves as a binary object mask that highlights salient contours. Next, to extract the foreground object, we perform element-wise multiplication:
(3) $I_{f}=I_{r}\odot I_{b}$
@@
where $\odot$ denotes element-wise multiplication, and yielding $I_{f}$, the foreground object image with the background suppressed.
@@
Step 2: Image CLIP Embedding Processing. To capture semantic information from each image representation, we utilize a frozen CLIP-ViT-H/14 to extract embeddings as follows:
(4) $C_{b}=\text{CLIP}(I_{b}),\quad C_{f}=\text{CLIP}(I_{f}),\quad C_{r}=\text{CLIP}(I_{r})$
@@
where $C_{b}$, $C_{f}$, and $C_{r}$ represent the embeddings corresponding to the binary mask, the foreground object, and the raw scene image, respectively. These embeddings serve as the foundation for subsequent hierarchical contrastive learning with EEG features.
4.3. Hierarchical EEG encoding

### v3

# ViEEG: Hierarchical Visual Neural Representation for EEG Brain Decoding

@@ Abstract
Understanding and decoding brain activity into visual representations is a fundamental challenge at the intersection of neuroscience and artificial intelligence. While EEG visual decoding has shown promise due to its non-invasive, and low-cost nature, existing methods suffer from Hierarchical Neural Encoding Neglect (HNEN)—a critical limitation where flat neural representations fail to model the brain’s hierarchical visual processing hierarchy. Inspired by the hierarchical organization of visual cortex, we propose ViEEG, a neuro-inspired framework that addresses HNEN. ViEEG decomposes each visual stimulus into three biologically aligned components—contour, foreground object, and contextual scene—serving as anchors for a three-stream EEG encoder. These EEG features are progressively integrated via cross-attention routing, simulating cortical information flow from low-level to high-level vision. We further adopt hierarchical contrastive learning for EEG-CLIP representation alignment, enabling zero-shot object recognition. Extensive experiments on the THINGS-EEG dataset demonstrate that ViEEG significantly outperforms previous methods by a large margin in both subject-dependent and subject-independent settings. Results on the THINGS-MEG dataset further confirm ViEEG’s generalization to different neural modalities. Our framework not only advances the performance frontier but also sets a new paradigm for EEG brain decoding.
1 Introduction
@@
Electroencephalogram (EEG) visual decoding aims to bridge the gap between cortical dynamics and machine perception by reconstructing visual experiences from brain activity (Du et al. 2023). As a pivotal modality in brain-computer interfaces, EEG (Craik, He, and Contreras-Vidal 2019) has gained prominence due to its cost-effectiveness, portability, and millisecond-level temporal resolution. Recent years have witnessed substantial progress in EEG visual decoding methodologies (Liu et al. 2024; Bai et al. 2023; Fu et al. 2025). However, unlike functional magnetic resonance imaging (fMRI) signals (Worsley et al. 2002) that have established sophisticated decoding frameworks (Zafar et al. 2015; Sun et al. 2023; Horikawa and Kamitani 2017; Zhou et al. 2024), EEG visual decoding remains in infancy, with significant challenges unaddressed, particularly in modeling the brain hierarchical visual processing mechanisms.
Figure 1: Triple-view hierarchical visual decoding.
Figure 2: EEG brain visual decoding of ViEEG. In subgraph. B, red: ground-trues images; green: EEG reconstructed images..
@@
Contemporary EEG decoding approaches predominantly establish direct visual-semantic mappings between EEG and stimulus images. For instance, Mb2C (Wei et al. 2024) employed contrastive learning for direct EEG-image embedding alignment with contrastive language-image pretraining (CLIP) (Radford et al. 2021; Ramesh et al. 2022), while NICE (Song et al. 2024) developed spatiotemporal neural architectures tailored for EEG visual decoding. Recent innovations include incorporating brain topological connectivity (Li et al. 2024) and integrating multimodal semantic cues (Zhang et al. 2024). Despite these advancements, most existing methods overlook a fundamental property of the human visual system: its hierarchical nature. We term this issue as Hierarchical Neural Encoding Neglect (HNEN). Rooted in the canonical hierarchical cortical processing pathway, the human visual system processes stimuli through dissociable stages: from low-level edge detection to high-level semantic integration (Felleman and Van Essen 1991; Riesenhuber and Poggio 1999). In contrast, existing methods collapse biological hierarchy into flat representation learning, resulting in incomplete visual information capturing.
@@
Substantial neuroscientific evidencee (Ben-Yishai, Bar-Or, and Sompolinsky 1995; Li, Todo, and Tang 2022) supports a hierarchically organized cortical visual pathway for visual perception: The early visual cortex (V1111Primary visual cortex (V1): basic edge detection and orientation processing /V2222Secondary visual cortex (V2): local contours and complex shapes integration (Bridge et al. 2005)) specializes in elementary low-level feature extraction (e.g., edge contours), the intermediate ventral stream (V4333Visual Area 4 (V4): shape/color information processing/IT [^4] Inferotemporal Cortex (IT): object recognition and semantic categorization (Hansen, Kay, and Gallant 2007; Arcaro and Livingstone 2021)) processes object-level semantics, while higher-order association cortex correlates [^5] Association Cortex Correlates (AC): Contextual scene understanding integration (Grill-Spector and Malach 2004) integrate high-level contextual scene understanding. This hierarchical visual decoding enables humans to progressively distill visual information—from coarse contours to fine-grained semantics—through layered cortical interactions. However, existing EEG brain decoding methods adopt a monolithic processing paradigm that adopts a flat representation paradigm that collapses biological hierarchy. By attempting direct mappings between EEG and holistic image embeddings, these approaches suffer from HNEN, failing to distinguish:
@@
  • Contextual scene attributes (AC correlates)
The absence of cortically-aligned feature disentanglement fundamentally limits the capacity to localize visual saliency, particularly in preserving contour fidelity.
@@
  • A biologically-inspired hierarchical framework for EEG visual decoding, pioneering tri-stream modeling (Contour $\to$ Object $\to$ Context). To our knowledge, this is the first work to explicitly enforce feature disentanglement along the visual cortical hierarchy.
@@
  • New state-of-the-art performance on THINGS-EEG: achieving 40.9% Top-1 ($\uparrow$49.82%) in subject-dependent and 22.9% Top-1 ($\uparrow$45.86%) in cross-subject settings, substantially outperforming previous benchmarks.
2 Related Works
Figure 3: Overview of the ViEEG framework. The input image is decomposed into three biologically inspired views: binary mask, foreground object, and raw scene. Corresponding EEG responses are encoded by parallel spatiotemporal encoders, integrated via hierarchical cross-attention, and aligned with CLIP embeddings through hierarchical contrastive learning.
@@ 2 Related Works
Decoding the brain’s response to visual stimuli has been a central pursuit in neuroscience and brain-computer interface (BCI) research (Lee, Lee, and Lee 2020; Shi et al. 2024; Liu et al. 2025). Early efforts predominantly focused on fMRI visual decoding (Allen et al. 2022; Huo et al. 2024), where high spatial resolution inherently captures hierarchical visual processing. For instance, MindEye (Scotti et al. 2023), proposed to map fMRI to CLIP features via diffusion models, while MindBridge (Wang et al. 2024) developed to align fMRI with images and textual captions. These methods leverage fMRI’s spatial precision to distinguish low-level (e.g., edges) and high-level (e.g., semantics) visual features implicitly, even when relying on simplistic alignment.
@@
In contrast, the spatial ambiguity of the EEG centimeter scale forces existing methods (Rakhimberdina et al. 2021; Song et al. 2024) into flat representation paradigms that confuse hierarchical visual processing. As the dominant brain decoding approach of fMRI and EEG, direct EEG-image alignment fails to model the brain’s progressive refinement from edges to semantics, a limitation we term HNEN. While some attempts incorporate electrode topology (Li et al. 2024) or depth cues (Zhang et al. 2024), these merely address symptoms (e.g., spatial noise) rather than HNEN’s root cause: EEG’s inability to functionally isolate cortical processing stages. This divergence highlights a critical gap: fMRI passively inherits hierarchical decoding through anatomy, whereas EEG demands active hierarchy simulation. Our proposed ViEEG bridges the gap via artificial cortical anchors—decomposing images into low-level edge masks, foreground entities, and high-level contextual scenes. By enforcing hierarchical EEG embedding disentanglement, we circumvent EEG’s spatial limitations while aligning with biological visual processing—a novel paradigm transcending both fMRI’s anatomy-dependence and EEG’s flat representations.
3 Method
3.1 Preliminaries
@@
$E$ ${R}^{C\times T}$ Raw EEG signals
II ${R}^{H\times W}$ Raw image
$I_{b}$ ${R}^{D\times D}$ Binary object mask image
$I_{f}$ ${R}^{D\times D}$ Foreground object image
$I_{r}$ ${R}^{D\times D}$ Raw scene image
$F_{b},F_{f},F_{r}$ ${R}^{d}$ Contour/object/context EEG embedding
$C_{b},C_{f},C_{r}$ ${R}^{d}$ Contour/object/context CLIP embedding
@@ II ${R}^{H\times W}$ Raw image
Table 1: Key symbol definitions. $C$: EEG channels, $T$: timepoints, $H\times W$: image resolution, $D\times D$: reshaped image resolution (512), $d$: embedding dimension (1024).
@@
We formulate EEG visual decoding as a hierarchical zero-shot object recognition, where test categories are disjoint from the training set. Given EEG recordings $E$ with $N$ samples elicited by visual stimuli $\{I_{b},I_{f},I_{r}\}$, our objective is to learn hierarchical EEG embeddings $\{F_{b},F_{f},F_{r}\}$ that align with the corresponding CLIP embeddings $\{C_{b},C_{f},C_{r}\}$ through hierarchical visual decoding. The learning objective for zero-shot object recognition is formulated as:
$\min_{\theta}\frac{1}{N}\sum_{i=1}^{N}D(F^{i},C^{i}),$ (1)
@@
where $D(\cdot,\cdot)$ measures feature similarity. On the THINGS-EEG dataset, we train with class-disjoint stimuli and test on unseen categories. At inference time, the concatenated EEG embeddings $F_{b},F_{f},F_{r}$} are compared to CLIP embeddings using cosine similarity within an InfoNCE-style contrastive learning framework (Radford et al. 2021).
3.2 Overall Architecture
@@
ViEEG introduces a biologically inspired hierarchical architecture for EEG visual decoding, simulating layered cortical processing. As illustrated in Figure 2, ViEEG comprises three synergistic components: (1) Hierarchical image decomposition, (2) Hierarchical EEG encoding, and (3) Hierarchical Contrastive Learning. During initialization, we extract CLIP embeddings from the three decomposed visual representations. In the training phase, EEG features extracted from EEG encoder are aligned with the image CLIP embeddings using contrastive learning. During testing, EEG embeddings extracted from the test set are matched to target samples based on similarity measurements for evaluation.
3.3 Hierarchical Image Decomposition
@@
Inspired by the hierarchical structure of the human visual system, we decompose each stimulus image $I$ into three biologically aligned representations via semantic segmentation.
@@
Step 1: Image Processing. We employ the pre-trained BiRefNet [^6] https://huggingface.co/ZhengPeng7/BiRefNet (Zheng et al. 2024), a state-of-the-art high-resolution segmentation model, to generate hierarchical visual representations. The raw image $I$ is first normalized to yield $I_{r}$, ensuring consistent dimensions and intensity scaling. Saliency detection is then applied, and a binary contour representation is obtained via thresholding:
$I_{b}=I\left(\mathcal{F}_{\text{BiRefNet}}(I_{r})>\tau\right)$ (2)
@@
where $\mathcal{F}_{\text{BiRefNet}}$ denotes the pre-trained BiRefNet model, $I(\cdot)$ is the indicator function, and $\tau$ is threshold value. The resulting $I_{b}$ serves as a binary object mask that highlights salient contours. Next, to extract the foreground object, we perform element-wise multiplication:
$I_{f}=I_{r}\odot I_{b}$ (3)
@@
where $\odot$ is element-wise multiplication, and yielding $I_{f}$, the foreground object image with the background suppressed.
@@
Step 2: Image CLIP Embedding Processing. To capture semantic information from each image representation, we utilize a frozen CLIP-ViT-H/14 to extract embeddings as:
$C_{b}=\text{CLIP}(I_{b}),\quad C_{f}=\text{CLIP}(I_{f}),\quad C_{r}=\text{CLIP}(I_{r})$ (4)
@@
where $C_{b}$, $C_{f}$, and $C_{r}$ represent the embeddings corresponding to the binary mask, foreground object, and raw scene image, respectively. These embeddings serve as the foundation for subsequent hierarchical contrastive learning with EEG.
3.4 Hierarchical EEG Encoding